# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation in Low-Data Regimes

### v4

@@ 5 Results and Discussion
$\quad$-- proxy 85.9 67.1 51.0 88.0 87.4 61.3 63.5 70.8 72.6
$\quad$-- sonar 88.4 69.8 57.3 91.3 90.9 65.7 68.4 72.8 76.2
@@
$\quad$-- proxy 85.6 61.8 46.2 87.9 84.5 57.8 59.1 64.8 69.1
$\quad$-- sonar 82.1 58.8 45.5 86.0 85.5 53.3 57.2 63.3 67.3
Table 4: Results on the Casablanca dataset. Best results are shown in bold. Second best results are underlined. WER (↓) scores are reported after normalization and removing diacritics. We report the score on the test split of each dataset.
@@ 3.1 CER Results
Model Size CV15.0 MGB2 MGB3 MGB5 Fleurs In-house Data SADA
ALG JOR PAL UAE YEM
Baselines
Amazon -/- -/- -/- -/- -/- -/- 70.2 25.6 29.0 40.8 43.5 -/-
[…]
HuBERT 0.31 18.9 17.3 9.5 45.5 10.9 44.3 23.3 27.9 36.7 38.8 34.5
W-FT 1.5 21.9 8.1 26.9 62.3 3.4 69.6 37.2 35.4 69.1 64.8 65.7
MMS-all 1.0 80.9 13.4 34.6 45.9 6.3 78.0 55.4 75.1 78.1 76.6 38.0
[…]
SM4T-L-v1 2.3 7.3 10.5 22.6 52.1 5.1 47.8 18.8 23.1 27.4 32.5 37.8
SM4T-L-v2 2.3 3.5 8.7 18.6 53.7 4.0 52.0 14.6 17.2 23.3 30.7 41.8
[…]
W-M 0.77 13.2 18.5 39.5 88.3 5.1 69.9 21.1 24.7 52.6 52.0 74.1
W-L-v2 1.5 7.8 15.3 33.0 68.9 3.6 71.7 17.0 22.3 38.2 45.5 51.2
W-L-v3 1.5 5.2 7.6 17.3 44.6 3.2 65.4 16.3 22.7 32.7 38.9 45.6
DW-16-16 0.80 7.2 10.8 25.1 43.3 6.6 38.5 18.2 23.3 27.7 31.6 38.9
DW-32-16 1.12 5.9 8.9 21.4 40.4 4.8 33.4 14.7 19.5 22.8 28.1 49.2
[…]
DW-32-16++ 1.12 5.5 8.8 20.3 40.6 3.1 33.3 13.4 18.8 21.1 26.8 35.8
[…]
Ours
UDW-16-16 0.80
$\quad$-- nll 8.15 11.26 27.98 55.25 6.26 41.4 25.7 20.52 35.96 49.2 55.0
[…]
$\quad$-- pesq 7.5 10.6 26.4 51.0 5.3 43.2 17.2 22.7 30.9 36.4 48.1
$\quad$-- sonar 5.62 8.98 23.4 46.97 4.41 37 15.11 19.56 24.29 28.24 35.6
[…]
$\quad$-- proxy 4.8 10.4 24.3 48.0 4.8 41.6 16.3 21.1 27.5 33.0 35.3
$\quad$-- sonar 6.1 9.8 24.2 47.1 5.0 38.9 17.2 21.4 26.2 29.5 33.96
UDW-32-16-++ 1.12
$\quad$-- proxy 5.8 9.2 21.1 44.1 4.2 37.1 14.5 20.1 23.1 29.2 31.4
$\quad$-- sonar 5.3 9.9 22.4 44.2 3.9 34.6 19.2 15.1 23.2 27.0 33.4

### v5

@@ 5 Results and Discussion
$\quad$-- proxy 85.9 67.1 51.0 88.0 87.4 61.3 63.5 70.8 71.9
$\quad$-- sonar 88.4 69.8 57.3 91.3 90.9 65.7 68.4 72.8 75.6
@@
$\quad$-- proxy 85.6 61.8 46.2 87.9 84.5 57.8 59.1 64.8 68.5
$\quad$-- sonar 82.1 58.8 45.5 86.0 85.5 53.3 57.2 63.3 66.5
Table 4: Results on the Casablanca dataset. Best results are shown in bold. Second-best results are underlined. WER (↓) scores are reported after normalization and removing diacritics. We report the score on the test split of each dataset.
@@ 2.1 SADA Dataset
Table 7 summarizes the statistics of the SADA dataset used in our experiments.
Dialect Test(S/D) Valid (S/D)
Najdi 1703/2.0709 2249/3.3155
MTOS 1320/4.8044 1048/3.82
Khaliji 1150/1.1308 679/0.6317
Hijazi 809/1.1202 528/0.6423
Unknown 762/0.8325 489/0.4861
NA 167/0.1341 2/0.0004
MSA 157/0.5406 54/0.1682
Egyptian 96/0.0865 45/0.0524
Shamali 18/0.0243 -
Yemeni 7/0.0052 23/0.0349
Levantine - 19/0.0137
Total 6189/10.75 5136/9.17
@@ 3.1 CER Results
We report the character error rates (CER) across different settings and datasets in Table 3.1.
@@
Model Size CV15.0 MGB2 MGB3 MGB5 Fleurs In-house Data SADA
ALG JOR PAL UAE YEM
Baselines
Amazon -/- -/- -/- -/- -/- -/- 70.2 25.6 29.0 40.8 43.5 -/-
[…]
HuBERT 0.31 18.9 17.3 9.5 45.5 10.9 44.3 23.3 27.9 36.7 38.8 34.5
W-FT 1.5 21.9 8.1 26.9 62.3 3.4 69.6 37.2 35.4 69.1 64.8 65.7
MMS-all 1.0 80.9 13.4 34.6 45.9 6.3 78.0 55.4 75.1 78.1 76.6 38.0
[…]
SM4T-L-v1 2.3 7.3 10.5 22.6 52.1 5.1 47.8 18.8 23.1 27.4 32.5 37.8
SM4T-L-v2 2.3 3.5 8.7 18.6 53.7 4.0 52.0 14.6 17.2 23.3 30.7 41.8
[…]
W-M 0.77 13.2 18.5 39.5 88.3 5.1 69.9 21.1 24.7 52.6 52.0 74.1
W-L-v2 1.5 7.8 15.3 33.0 68.9 3.6 71.7 17.0 22.3 38.2 45.5 51.2
W-L-v3 1.5 5.2 7.6 17.3 44.6 3.2 65.4 16.3 22.7 32.7 38.9 45.6
DW-16-16 0.80 7.2 10.8 25.1 43.3 6.6 38.5 18.2 23.3 27.7 31.6 38.9
DW-32-16 1.12 5.9 8.9 21.4 40.4 4.8 33.4 14.7 19.5 22.8 28.1 47.3
[…]
DW-32-16++ 1.12 5.5 8.8 20.3 40.6 3.1 33.3 13.4 18.8 21.1 26.8 35.8
[…]
Ours
UDW-16-16 0.80
$\quad$-- nll 8.15 11.26 27.98 55.25 6.26 41.4 25.7 20.52 35.96 49.2 55.0
[…]
$\quad$-- pesq 7.5 10.6 26.4 51.0 5.3 43.2 17.2 22.7 30.9 36.4 48.1
$\quad$-- sonar 5.62 8.98 23.4 46.97 4.41 37 15.11 19.56 24.29 28.24 35.6
[…]
$\quad$-- proxy 4.8 10.4 24.3 48.0 4.8 41.6 16.3 21.1 27.5 33.0 35.3
$\quad$-- sonar 6.1 9.8 24.2 47.1 5.0 38.9 17.2 21.4 26.2 29.5 33.96
UDW-32-16-++ 1.12
$\quad$-- proxy 5.8 9.2 21.1 44.1 4.2 37.1 14.5 20.1 23.1 29.2 31.4
$\quad$-- sonar 5.3 9.9 22.4 44.2 3.9 34.6 19.2 15.1 23.2 27.0 33.4
@@ 3.2 Training Parameters
Table 14 lists the hyperparameters used for training our models across all experiments.
@@ 3.3 Results
We present additional experimental results evaluating orthographic variants in Table 15.